OpenStudy (anonymous):

The quadratic 2x2 + 6x + 3 has what type of roots

OpenStudy (hoblos):

Δ = 6^2 -4(2)(3) = 36 - 24 =12 > 0 so this quadratic has 2 real roots

OpenStudy (saifoo.khan):

Use discriminant. b^2 - 4ac

OpenStudy (angela210793):

how did u do ''delta'' @hoblos

OpenStudy (saifoo.khan):

crtl c + ctrl v

OpenStudy (saifoo.khan):

Δpple

OpenStudy (angela210793):

Very smart Saif :P

OpenStudy (hoblos):

well i dont really!! i have a copy :P

OpenStudy (angela210793):

ohhhhhh...i thought there was a "alt+_____" :D

OpenStudy (saifoo.khan):

hehe. ;)

OpenStudy (hoblos):

googled it..... alt+0916

OpenStudy (angela210793):

:D Thanks xD

OpenStudy (hoblos):

yw =)